OpenStudy (anonymous):

What is 1+1 divided by 1

OpenStudy (theprofessor35):

1 + 1 = what

OpenStudy (theprofessor35):

1 + 1 = 2

OpenStudy (isaidavila):

Is (1+1)/2 or 1+(1/2)?

OpenStudy (anonymous):

I'm in second grade

OpenStudy (theprofessor35):

anything divided by 1 = itself. 1+1 = 2 2/1 = 2

OpenStudy (anonymous):

Ok

OpenStudy (tennis5518):

If your in second grade you shouldn't be on this website!

alones (alones):

@Handgrip1 you're srsly a second grader

OpenStudy (mustbethestars):

To be honest you're too young for this website if you are truly in second grade